Linear Algebra: Let A = [{9,-6},{12,-8}] find A^100
Answer #1

A = [ 12 12 -8 Characters the equation af A 8 G-^)(- 8-1 +72 = 0 22_x=0 - X=0, X2 = 1 Let [x] be vector Wor. to the eigen 112A = 1 [3:09 ws لها بم 4 and A PDP-1 100 Aloo PD P-1 14 3. 3. -2 A loo ا - 1341 LoiJl 132113) A 100 - - 6 A 100 - 8 12